# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

### v1

@@ IV Exploring Parameter Selections
The choice of parameters $s$ (neighborhood size) and $\delta$ (neighborhood similarity threshold) affect the resulting clustering differently. The neighborhood size $s$ directly affects the defined neighborhoods for each observation and hence impacts the construction of the set $\mathbf{C}$ of filtered centers. Under the max strategy, for any given $s$, the value of $\delta$ further influences how the filtered centers are grouped. Basically, a smaller value of $\delta$ results in fewer, and larger, groups of filtered centers.
IV-A Hierarchical Structure of grouping under the max strategy
@@
Under the max strategy, given a certain $s$, all possible values of the number of groups can be derived. By starting at $\delta=\delta_{0}$ and progressively reducing $\delta$, all $\delta$’s at the time that the number of groups changes can be recorded.
@@
We use a toy example to illustrate the process, shown in Figure 5. Each square matrix is referred as a “grouping matrix”. The order of the matrix denotes the number of groups, and its entries identify the grouping critical points among these groups for a given $\delta$ value. In each step, the grouping matrix obtained from the previous one by merging columns and rows that contain the largest non-trivial (not 00 or $1$) value. The merging process retains the larger of the values in the corresponding entries of the columns and rows. Eventually, we are left with completely disjoint groups, represented by the identity matrix as the grouping matrix.
@@
The process of grouping generates two key lists: one for the number of groups and another for the corresponding threshold values. More precisely, in the toy example, the potential numbers of groups are $6$, $4$, $3$, and $2$, with the $\delta$ values ranging between the intervals $[0.6,\delta_{U})$, $[0.5,0.6)$, $[0.4,0.5)$, and $[0,0.4)$ at each stage. The hierarchical structure ensures that if the number of clusters is known or can be estimated using any methods (refer to [48] for estimation techniques), a suitable $\delta$ can be directly determined from the relationship between the lists $\boldsymbol{G}$ and $\boldsymbol{\delta}$.
IV-B Depth-Based Clustering Metric
The next challenge of parameter selection is to identify the optimal pair of parameters $(s,\delta)$, which requires an appropriate internal metric to determine the best clustering among various results. However, the DLCC under the max strategy often addresses clustering problems involving non-convex shapes or closely positioned clusters of significantly different sizes, where well-known internal metrics such as ASW and CHI encounter difficulties. The limitations of these metrics are illustrated using two synthetic datasets, SmileyF and Bainba — see Figures 1 and 6, and Table II. In both figures, the clustering results in (a) are the actual labels for the datasets, for which both ASW and CHI yield low scores. Both criteria awarded high scores to the clustering result (c) in the SmileyF dataset, although human assessment considers (c) to perform the poorest. Similarly, both criteria favour clustering result (b) for the Bainba dataset, which contains only two clusters. This discrepancy highlights the need for a metric that better handles clustering of complex shapes.
Figure 6: Clustering results for the synthetic dataset Bainba. Panel (a) shows the ground truth, (b) merges two distinct clusters into one, and (c) incorrectly clusters a point at the top of the central elliptical cluster.
TABLE II: ASW and CHI values corresponding to the clustering results in Figures 6 and 1.
DATA SET Label (a) (b) (c)
Bainba: ASW $0.341$ $0.491$ $0.339$
Bainba: CHI $642.13$ $842.52$ $641.99$
SmileyF: ASW $-0.146$ $-0.152$ $0.435$
SmileyF: CHI $45.25$ $43.18$ $1470.62$
The DC metric is introduced to evaluate and compare different clustering outcomes when the ground truth is not available, facilitating the identification of the optimal pair of parameters. The main idea is to calculate within- and between-cluster similarities in a small, fixed region instead of computing similarities globally to handle non-convex clusters. This approach, inspired by the DBCA methodology [33], is adapted in our context with a broader scope. While DBCA specifies similarity as $D_{\mathrm{MD}}^{i}(\mathbf{z})$, we generalize this concept to accommodate various forms of similarity matrices, $\mathbb{S}$. Some definitions from DBCA are listed below:
Definition 6 (Core Neighbors).
Given a depth similarity threshold $\eta$, a point $\mathbf{x}_{j}$ is a core neighbor of $\mathbf{x}_{i}$ (with respect to $\eta$) if $\mathbb{S}_{ij}\geq\eta$.
Definition 7 (Depth-Connection).
A point $\mathbf{x}_{u}$ is depth-connected to $\mathbf{x}_{v}$ if there is a chain of points $\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{i}$, where each two adjacent points are core neighbors for each other, $\tilde{\mathbf{x}}_{1}=\mathbf{x}_{u}$, and $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{v}$.
Definition 8 (DBCA-Cluster).
A chain $\mathbf{V}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{i}\}$ is called a maximal depth-connected chain if there is no $\mathbf{x}_{j}\notin\mathbf{V}$ satisfying $\sup_{\mathbf{x}_{i}\in\mathbf{V}}\mathbb{S}_{ij}\geq\eta$. Let $\tilde{n}$ be the minimum number of points required to form a cluster in DBCA, then each cluster is formed by observations that are in the same maximal depth-connected chain with length at least $\tilde{n}$.
Definition 9 (Outliers).
All observations that remain unlabelled in DBCA are regarded as outliers.
In Definition 6, larger $\eta$ corresponds to fewer core neighbors, which corresponds to a smaller fixed region. Let $\eta_{X}$ be the largest value of $\eta$ such that all observations in the dataset are depth-connected. Let $k$ be the label of a cluster in the clustering result which is to be examined by the DC metric, and $\eta_{k}$ be the maximum $\eta$ value such that the cluster is decomposed into a single DBCA-cluster and outliers (without considering observations in other clusters). Let $\mathcal{X}_{k}$ be the set of observations in cluster $k$. Let $n_{k}=|\mathcal{X}_{k}|$ and $\mathbf{O}_{k}$ be the set of DBCA-outliers. For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, we define $\mathcal{J}_{k}^{i}=\{j|\mathbf{x}_{j}\in\mathcal{X}_{k},\mathbf{x}_{j}\neq\mathbf{x}_{i},\mathbb{S}_{ij}>\eta_{k}\}$ as the set of indices. Following this, the within cluster similarity $J_{k}$ of cluster $k$ is defined by
$J_{k}=\sum_{i=1}^{n_{k}}w_{i}\mu_{i},$
where
$\mu_{i}$ $=$ $\begin{cases}\dfrac{\sum\limits_{j\in\mathcal{J}_{k}^{i}}\mathbb{S}_{ij}}{|\mathcal{J}_{k}^{i}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{O}_{k},\\ \max\{\mathbb{S}_{ij}\mid\mathbf{x}_{j}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}\},\text{ if }\mathbf{x}_{i}\in\mathbf{O}_{k},\end{cases}$
$w_{i}$ $=$ $\begin{cases}\dfrac{|\mathcal{J}_{k}^{i}|}{\sum\limits_{\mathbf{x}_{v}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}}|\mathcal{J}_{k}^{v}|+|\mathbf{O}_{k}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{O}_{k},\\ \dfrac{1}{\sum\limits_{\mathbf{x}_{v}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}}|\mathcal{J}_{k}^{v}|+|\mathbf{O}_{k}|},\text{ if }\mathbf{x}_{i}\in\mathbf{O}_{k}.\end{cases}$
The weights $w_{i}$ in $J_{k}$ give observations with more core neighbors more weight and give outliers less weight. By considering DBCA-outliers, the DC metric can be made more robust by avoiding the significant adverse impact of an individual outlier.
For the cluster $k$, the set of inner observations is
$\mathbf{I}_{k}:=\left\{\mathbf{x}_{i}\in{\mathcal{X}_{k}}:\mathbb{S}_{ij}<\eta_{X}\text{ for all }\mathbf{x}_{j}\in{\mathbf{X}\setminus\mathcal{X}_{k}}\right\},$
where $\mathbf{X}$ is the total dataset. If $\mathbf{x}_{i}\in\mathbf{I}_{k}$ then $\mathbf{x}_{i}$ is called an inner observation of cluster $k$, i.e., observations in $\mathcal{X}_{k}$ that are not similar to any observation outside of $\mathcal{X}_{k}$. For defining the between cluster similarity $H_{k}$ of cluster $k$, we set $\mathcal{H}_{k}^{i}=\{h|\mathbf{x}_{h}\notin\mathcal{X}_{k},\mathbb{S}_{ih}>\eta_{X}\}$ to include indices of observations from other clusters that are core neighbors of points in $\mathcal{X}_{k}$ when $\eta=\eta_{X}$. $H_{k}$ is then defined as
$H_{k}=\sum_{i=1}^{n_{k}}\tilde{w}_{i}\tilde{\mu}_{i},$
where,
$\tilde{\mu}_{i}$ $=$ $\begin{cases}\dfrac{\sum\limits_{h\in\mathcal{H}_{k}^{i}}\mathbb{S}_{ih}}{|\mathcal{H}_{k}^{i}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{I}_{k},\\ \eta_{X},\text{ if }\mathbf{x}_{i}\in\mathbf{I}_{k},\end{cases}$
$\tilde{w}_{i}$ $=$ $\begin{cases}\dfrac{|\mathcal{H}_{k}^{i}|}{\sum\limits_{\mathbf{x}_{u}\in\mathcal{X}_{k}\setminus\mathbf{I}_{k}}|\mathcal{H}_{k}^{u}|+|\mathbf{I}_{k}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{I}_{k},\\ \dfrac{1}{\sum\limits_{\mathbf{x}_{u}\in\mathcal{X}_{k}\setminus\mathbf{I}_{k}}|\mathcal{H}_{k}^{u}|+|\mathbf{I}_{k}|}\,\text{if }\mathbf{x}_{i}\in\mathbf{I}_{k}.\end{cases}$
In $H_{k}$, the weights $\tilde{w}_{i}$ give observations which are close to other clusters more weight, but less weight for inner observations.
The total DC metric for a clustering result is then given by
$\text{DC}=\sum_{k}\frac{n_{k}}{n}(J_{k}-H_{k}),$ (20)
which gives more weight to the clustering performance in clusters with more observations. Some properties of the DC metric are summarized as follows:
  1. Property 1:
    For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, reducing the threshold from $\eta_{k}$ to $\eta_{k^{*}}$ implies $\mu_{i^{*}}\leq\mu_{i}$.
  2. Property 2:
    For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, reducing the threshold from $\eta_{k}$ to $\eta_{k^{*}}$ implies $\tilde{\mu}_{i^{*}}\leq\tilde{\mu}_{i}$.
  3. Property 3:
    For any cluster $k$, the $\mu_{i}$ given by any outlier is less than $\eta_{k}$.
  4. Property 4:
    For any cluster $k$, $H_{k}\geq\eta_{X}$, higher proportions of inner observations in $k$ leads to smaller $H_{k}$.
Table III displays the DC metrics for the aforementioned clustering results for the Bainba and SmileyF datasets. Notably, the ranking of these scores aligns more closely with human assessments. Thus, using the DC metric allows for the selection of the parameters for the DLCC algorithm under the max strategy.
TABLE III: The DC metrics for clustering results in Figures 6 and 1, where $\tilde{n}=3$. Note this score correctly assesses (a) as the best assignment of labels among the three panels.
(a) (b) (c)
Bainba $0.046$ $0.044$ $0.045$
SmileyF $0.174$ $0.164$ $0.079$
@@ V Experiments and Evaluations
In this section, we evaluate the performance of DLCC under various strategies, comparing it with several other clustering algorithms on both synthetic and real-world datasets.

### v2

@@ IV-B Max Strategy
The choice of the parameters $s$ (neighborhood size) and $\delta$ (neighborhood similarity threshold) affect the resulting clustering differently. The neighborhood size $s$ directly affects the defined neighborhoods for each observation and hence impacts the construction of the set $\mathbf{C}$ of filtered centers. Under the max strategy, for any given $s$, the value of $\delta$ further influences how the filtered centers are grouped. Basically, a smaller value of $\delta$ results in fewer, and larger, groups of filtered centers.
IV-B1 Hierarchical Structure of Grouping
@@
Given a certain $s$, all possible values of the number of groups can be derived. By starting at $\delta=\delta_{0}$ and progressively reducing $\delta$, all $\delta$’s at the time that the number of groups changes can be recorded.
@@
We use a toy example to illustrate the process, shown in Figure 5. Each square matrix is referred as a “grouping matrix”. The order of the matrix denotes the number of groups, and its entries identify the grouping critical points among these groups for a given $\delta$ value. In each step, the grouping matrix is obtained from the previous one by merging columns and rows that contain the largest non-trivial (not 0 or $1$) value. The merging process retains the larger of the values in the corresponding entries of the columns and rows. Eventually, we are left with completely disjoint groups, represented by the identity matrix as the grouping matrix.
@@
The process of grouping generates two key lists: one for the number of groups and another for the corresponding threshold values. More precisely, in the toy example, the potential numbers of groups are $6$, 44, 33, and $2$, with the $\delta$ values ranging between the intervals $0.6,\delta_{U})$, $[0.5,0.6)$, $[0.4,0.5)$, and $[0,0.4)$ at each stage. The hierarchical structure ensures that if the number of clusters is known or can be estimated using any methods, a suitable $\delta$ can be directly determined from the relationship between the lists $\boldsymbol{G}$ and $\boldsymbol{\delta}$.
IV-B2 Parameter selection
For the max strategy, clusters are formed by connecting multiple local convex regions. Thus, the neighborhood size $s$ should be relatively small. In our experiments, we recommend $s\in[20,60]$. For each $s$, we generate a corresponding candidate list of thresholds $\boldsymbol{\delta}$ and evaluate the resulting solutions accordingly. Note that the construction of the similarity matrix is independent of the parameters. Therefore, it can be computed once in advance before the parameter selection stage.
Without access to ground truth, clustering is inherently exploratory. As noted by Hennig [49], there is in general no unique “true” clustering, and validation often rely on a combination of internal validation indices (IVI) and subjective assessment, where visualization plays an important supporting role. Our parameter selection rule follow this philosophy.
In terms of IVIs, classical criteria such as the silhouette width [50] and the Calinski-Harabasz index [51] are primarily designed for compact and approximately convex clusters, and are not suitable for clusters with arbitrary shapes. To better reflect arbitrary geometries, we adopt the cluster validity index based on density-involved distance (CVDD) [52]. [^1] https://github.com/hulianyu/CVDD CVDD replaces the usual notion of separation by a density-aware connectivity distance defined on a kkNN graph. Two points are regarded as close only if they can be connected by a path that stays in relatively high-density regions. Consequently, CVDD penalizes clustering results that effectively require “crossing” low-density gaps to maintain within-cluster connectivity, while remaining insensitive to non-convex shapes as long as they are density-connected. The resulting score balances density-aware between-cluster separation against within-cluster connectivity; further details are referred to Hu and Zhong [52].
However, IVIs inevitably encode particular assumptions about “separation” and “cohesion”, and tuning based on a single IVI alone is therefore often insufficient in practice. Indeed, Lewis et al. [53] reported that human assessments can be more consistent than existing IVIs when judging clusterings. We therefore combine CVDD with visualization for model selection. Specifically, we run DLCC over the grid of $(s,\boldsymbol{\delta})$, which yields solutions with varying estimated numbers of clusters $\hat{K}$. For each distinct $\hat{K}$, we retain the solution that attains the largest CVDD value as its representative. We then visualize the representatives of the top three $\hat{K}$ values ranked by CVDD and choose the final model by visual assessments.
In experiments shown in Figure 7, although CVDD identifies suitable parameter pairs for most datasets, it may not always favor the most appropriate $\hat{K}$. For instance, on the Bainba dataset (LABEL:subfig:bainba), CVDD actually prefers merging the two upper groups due to their proximity and similar local densities; visually, however, they form two distinct clusters and should remain separated.
IV-C Min strategy
Figure 6: Empirical association between the proposed depth-connectivity score (blue; left axis) and the ARI (orange; right axis) as the neighborhood size $s$ varies under the min strategy with a fixed $K$.
Under the min strategy, particularly for real-world high-dimensional data (e.g., $d>50$), selecting parameters solely using a single IVI is often challenging. Nevertheless, IVIs (e.g., CVDD) remain useful for narrowing down reasonable ranges of $s$ and estimate $\hat{K}$. With a known or an estimated $K$, we recommend choose $s\in[0.6\,n/K,\,n/K]$. To further evaluate candidate clusterings, we then exploit the symmetrized depth-based similarity matrix $\mathbb{S}$. Before proceeding, we introduce the reachable similarity, a max-min (bottleneck-path) connectivity measure in graph theory [54, 55].
Definition 6 (Reachable similarity).
Let $S$ be a symmetric similarity matrix with entries in $[0,1]$. The reachable similarity between nodes $i$ and $j$ is
$\mathrm{rs}(i,j)=\max_{\boldsymbol{p}:i\to j}\ \min_{(u,v)\in\boldsymbol{p}}S_{u,v},$
i.e., the strongest bottleneck similarity over all paths connecting $i$ and $j$.
Given a candidate partition $\{\pi_{g}\}_{g=1}^{K}$, we quantify connectivity using two pointwise quantities: the intra-cluster reachable similarity (IR) and the between-cluster reachable similarity (BR). For each point with label $\widehat{y}_{i}$, its IR is defined as the average reachable similarity from $\mathbf{x}_{i}$ to other points within the same cluster, where $\mathrm{rs}(\cdot,\cdot)$ is computed on the subgraph induced by nodes in $\pi_{\widehat{y}_{i}}$.
To define BR, we first identify the nearest competing cluster for $\mathbf{x}_{i}$ by comparing $\mathbf{x}_{i}$ to filtered centers from other clusters
$\tilde{g}_{i}=\arg\max_{g\neq\widehat{y}_{i}}\ \max_{\mathbf{x}_{u}\in\mathbf{g}_{g}}\ \mathbb{S}_{i,u},$
where $\mathbf{g}_{g}$ is the $g$th group of filtered centers. We then set BRi to be the average reachable similarity from $\mathbf{x}_{i}$ to points in the competing cluster $\pi_{\tilde{g}_{i}}$, with $\mathrm{rs}(\cdot,\cdot)$ computed on the subgraph induced by nodes in $\pi_{\tilde{g}_{i}}$.
Finally, we summarize a clustering result by the mean gap
$\texttt{Connectivity}=\frac{1}{n}\sum_{i=1}^{n}\big(\texttt{IR}_{i}-\texttt{BR}_{i}\big),$ (19)
where larger values indicate stronger within-cluster connectivity relative to the most competitive cross-cluster connectivity. We observe that when $K$ is fixed and ground truth labels are available, this depth-connectivity score exhibits a strong positive association with external agreement measures such as the adjusted Rand index (ARI; [56]) across a range of datasets; representative examples are shown in Fig. 6. This observation motivates using (19) as a practical tool for ranking candidate neighborhood sizes $s$ under the min strategy when ground truth labels are unavailable.
@@ V Experiments and Evaluations
In this section, we compare DLCC under both strategies with several competing clustering methods on synthetic and real-world data sets. In DLCC, the final-stage classifier is selected via CVDD; when RF is chosen, we report the median ARI, NMI, and CE over $100$ independent runs to account for its stochasticity. Detailed records of the selected classifiers are provided in the supplementary material.
V-A External Metric